The solubility of silver chloride at 25° C equals 1,34·10-4mol/l. Calculate the solubility product of this salt

Grade:12

1 Answers

FAIZANUL MUKHTAR
4 years ago
AgCl being a binary salt.
 
The relation between solubility product and solubility for binary salt is as:
       Ksp = S ×S
       Ksp= (1.34×10-4) (1.34×10-4)
             =1.8×10-8 mol2/L2